# RADE: A Neural Codec for Transmitting Speech over HF Radio Channels

### v1

@@ 1 Introduction
High-frequency (HF) push-to-talk (PTT) radio has the benefit of operating without infrastructure over ranges of several thousands of km. Applications include humanitarian, remote area, emergency and government communication when access to cellular and satellite systems cannot be guaranteed. The HF radio signal propagates from the transmitter to the receiver via reflection from upper layers of the atmosphere. Typically the signal is reflected multiple times by various sub-layers thus multiple, time shifted versions of the signal arrive and are summed at the receiver (multipath propagation). Single side band (SSB) – an analog communications system that was invented in 1915 [1] – entered widespread use in the 1950’s [2] and remains the de facto standard on HF due to its power and bandwidth efficiency, and robustness to multipath propagation. The voice quality of analog and digital HF speech systems remains low compared to modern cellular and Internet telephony services and has not seen significant improvement in 70 years. Key requirements for HF voice services are voice quality, narrow RF bandwidth, low SNR operation, robustness to multipath channels, and latency of less than 200 ms to support PTT speech.
@@
Consider the classical DSP speech over radio system in Fig. 1. Speech samples from a microphone are compressed to a low bit rate using a speech encoder. Forward error correction (FEC) adds redundant bits to protect the sensitive payload speech bits from channel errors. The bitstream is then converted to a signal suitable for transmission over the radio channel using a modulator (e.g. a sequence of QAM symbols mapped to OFDM sub-carriers). The received radio signal is demodulated to a bit stream; the FEC decoder attempts to correct any bit errors, and the speech decoder converts the signal back into a sampled speech signal for replay over a loudspeaker. Classical DSP digital voice systems have the following drawbacks: separating the speech coding and channel protection (FEC) leads to inefficiencies; the use of largely linear DSP means the inability to exploit non-linear dependencies when compressing voice; difficulty in handling multipath channels; long latencies due to use of interleavers/FEC to overcome multipath fading; and low quality speech from low bit rate classical DSP vocoders. They exhibit an abrupt threshold SNR where the system ceases to work, and speech quality does not gracefully scale with available channel SNR (although step changes are possible with mode switching).
@@
The use of ML combines the operations of quantisation, channel encoding and modulation, and allows joint training of the entire end-to-end system to minimise distortion when subjected to the HF channel impairments [5, 6]. The powerful non-linear transforms and prediction available in ML allow us to more efficiently model the speech signal and time based evolution of the channel to improve robustness. We employ the FARGAN vocoder [7] for high quality neural speech synthesis, however our work is applicable to any neural and even classical vocoders with a similar feature set.
Our contributions in this paper are:
@@
-50Imag050100-100-1000Real0.20.4
@@
Figure 3: 2D histogram of complex RADE encoder QAM symbols with all elements of $\mathbf{z}$ superimposed for a 60 second sample containing multiple speakers. The bin count has been scaled to a maximum of 1. Compared to digital QAM constellations, the RADE constellation looks like noise. Plotting individual elements of $\mathbf{z}$ produces a similar histogram (no obvious structure).
@@ 2 RADE Design
    We can perform acquisition, synchronisation, and sample rate conversion using well known DSP techniques, further simplifying the ML processing.
@@
The OFDM frame is shown in Fig. 4. Pilot symbols are periodically inserted into each OFDM carrier. After the IDFT stage, a cyclic prefix is inserted to guard against inter-symbol interference. To achieve an efficient ratio of pilots to data symbols, we place the QAM symbols from three consecutive $\mathbf{z}$ vectors (120 complex QAM symbols) in each OFDM frame, leading to an OFDM frame duration (and algorithmic delay) of 120 ms.
At the receiver the pilots are used to estimate the time-varying phase of the channel (equalisation), and for initial acquisition (coarse frequency, frame sync) of the received signal. Phase equalisation also allows small frequency offsets (±$\pm 2$ Hz) to be handled and frequency drift tracked. Neural networks are sensitive to magnitude scaling, so we also use the pilot symbols for coarse magnitude equalisation (gain control).
The insertion of pilot symbols and the cyclic prefix consume carrier power that would otherwise be available for payload symbols, and require the symbol rate and hence overall RF bandwidth to be increased to maintain the payload symbol rate. The overheads for RADE total 4 dB, including a 2-dB difference from ideal performance in the least squares estimation algorithm used for phase equalisation. The resulting waveform has a RF bandwidth of approximately 1500 Hz, with 500 Hz due to overheads.
@@ 3 Training
which simulates the saturation of a transmitter power amplifier by compressing the magnitude but retaining the phase. We reasoned that a bottleneck applied to the magnitude of the complex time domain signal would encourage the network to maximise the RMS power (and hence minimse the PAPR) given the channel noise and peak power constraint of the bottleneck.
To train we need to apply the $\mathrm{ctanh}(x)$ bottleneck in the time domain, and simultaneously apply a multipath channel model. Applying the multipath model in the time domain introduces phase rotations and inter-symbol interference (ISI), which would then require equalisation and removal inside the training loop. While possible this would require significantly slow down the training process, and require training at the higher rate $F_{s}$ sample rate ($F_{s}/R_{s}=160$ in our design).
@@ 4 Evaluation and Results
Informal listening tests on simulated and over the air samples suggest RADE significantly outperforms SSB. Since intelligibility – more than quality – is the primary goal for HF radio, we use Automatic Speech Recognition (ASR) to evaluate the performance of the proposed system. Five hundred samples from the Librispeech dataset were passed through RADE, SSB, and FreeDV 700D simulations at a range of SNRs, then post processed by the Whisper ASR system [12], and the Word Error Rate (WER) measured (Fig. 6). SSB was simulated by band limiting the input speech to 300-2700 Hz, and applying Hilbert compression such that the mean PAPR was around 8 dB. FreeDV 700D is an open-source HF digital voice protocol using an OFDM modem, rate $1/2$ FEC, and the Codec 2 classical vocoder [13]. The MPP channel was simulated using the time domain Watterson model Eq. (4). The Librispeech speech and Watterson model $G_{1}$ and $G_{2}$ datasets used for the evaluation were not part of the training dataset.
We consider two thresholds (a) Link closure - the point where barely intelligible speech can be sent over the system and (b) Good quality, effortless communication. At the 30% WER level (link closure), the ASR results indicate a 4 dB improvement for RADE over SSB for both AWGN and MPP channels. At the 5% WER level (good quality), the improvement is 13 dB for both channels. FreeDV 700D exhibits low speech quality with the Librispeech dataset, although we note it is in common use on HF by trained operators. Similarly, skilled operators can use SSB down to 0 dB SNR. Note the sharp knee in the FreeDV 700D AWGN curve at $-2$ dB SNR, common in digital speech system due to the abrupt breakdown of the FEC. RADE and SSB have a more desirable gradual trade off between SNR and speech quality.
@@
Results using FARGAN on clean features show that the resulting WER is very close to that of the clean speech. The use of our classical features and the choice of vocoder are thus not a limiting factor in the performance of the proposed system.
@@
Figure 6: Word error rate % versus SNR for simulated AWGN and MPP channels, with clean and uncoded FARGAN as controls. SNR for all signals is normalised to a 3000 Hz noise bandwidth.
@@ 4.1 Complexity
The encoder and decoder each require 1 MB of read-only storage (1M weights with 8-bit quantization) and require 32 MMACs for real-time operation. Both easily operate in real time on a laptop using an unoptimised PyTorch implementation. The overall complexity of the complete system is dominated by the FARGAN vocoder’s 300 MMACs [7].
@@ 4.2 Over the Air Demonstration
To test RADE over real world HF channels licensed Amateur Radio operators from around the world were invited to record a 10 second input sample of their own voice. This was converted to a RADE waveform sample, and concatenated with a Hilbert-compressed version of the same input sample to emulate an analog SSB signal with a 6-8 dB PAPR. Participants then played the concatenated sample through their SSB transmitters over real world HF Radio channels to remote KiwiSDR receivers of their choice (KiwiSDRs are SSB radio receivers that are connected to the public Internet). The received off air signal was processed to obtain a file of SSB and RADE output speech samples, and an estimate of channel SNR. The use of stored files enabled the same Tx signal to be transmitted to different KiwiSDR receivers under different channel conditions, and different transmit power levels. Tests were performed in English, Japanese, Cantonese, and German, over a variety of HF channels of up to 14,000 km (e.g. direct transmission from North America to Australia). Speech samples are available at [14].
@@ 5 Conclusion
We have combined a ML vocoder, ML autoencoder and classical DSP OFDM to build a system capable of sending speech over HF radio channels. It is robust to AWGN and multipath channel impairments, and the transmit signal has a PAPR of less than 1 dB. Unusually for HF speech systems, the audio bandwidth is 8000 Hz, despite requiring just 1500 Hz of RF bandwidth. Our ASR simulation and real world demonstration show performance significantly exceeding that of the analog SSB at the same SNR. Unlike classical DSP HF systems, speech quality improves gradually with channel SNR without any mode switching. Interestingly, our system also shows robustness to channel impairments we did not train for, e.g. impulse noise.

### v2

@@ 1 Introduction
High-frequency (HF) push-to-talk (PTT) radio has the benefit of operating without infrastructure over ranges of several thousands of km. Applications include humanitarian, remote area, emergency and government communication when access to cellular and satellite systems cannot be guaranteed. The HF radio signal propagates from the transmitter to the receiver via reflection from upper layers of the atmosphere. Typically the signal is reflected multiple times by various sub-layers thus multiple, time shifted versions of the signal arrive and are summed at the receiver (multipath propagation). Single side band (SSB) – an analog communications system that was invented in 1915 [1] – entered widespread use in the 1950’s [2] and remains the de facto standard on HF due to its power and bandwidth efficiency, and robustness to multipath propagation. The voice quality of analog and digital HF speech systems remains low compared to modern cellular and Internet telephony services and has not seen significant improvement in 70 years. Key requirements for HF voice services are voice quality, narrow RF bandwidth, low SNR operation, robustness to multipath channels, and one way latency of less than 200 ms [3] to support PTT speech.
@@
Consider the classical DSP speech over radio system in Fig. 1. Speech samples from a microphone are compressed to a low bit rate using a speech encoder. Forward error correction (FEC) adds redundant bits to protect the sensitive payload speech bits from channel errors. The bitstream is then converted to a signal suitable for transmission over the radio channel using a modulator (e.g. a sequence of QAM symbols mapped to OFDM sub-carriers). The received radio signal is demodulated to a bit stream; the FEC decoder attempts to correct any bit errors, and the speech decoder converts the signal back into a sampled speech signal for replay over a loudspeaker. Classical DSP digital voice systems have the following drawbacks: separating the speech coding and channel protection (FEC) leads to inefficiencies; the use of largely linear DSP means the inability to exploit non-linear dependencies when compressing voice; difficulty in handling multipath channels; long latencies due to use of interleavers/FEC to overcome multipath fading; and low quality speech from low bit rate classical DSP vocoders. They exhibit a threshold SNR where the system ceases to work, and speech quality does not gracefully scale with available channel SNR (although step changes are possible with mode switching).
@@
The use of ML combines the operations of quantisation, channel encoding and modulation, and allows joint training of the entire end-to-end system to minimise distortion when subjected to the HF channel impairments [6, 7, 8]. The powerful non-linear transforms and prediction available in ML allow us to more efficiently model the speech signal and time based evolution of the channel to improve robustness. We employ the FARGAN vocoder [9] for high quality neural speech synthesis, however our work is applicable to any neural and even classical vocoders with a similar feature set. Our contributions in this paper are:
@@
Refer to caption-50Imag050100-100-1000Real0.20.4
@@
Figure 3: 2D histogram of complex RADE encoder QAM symbols with all elements of $\mathbf{z}$ superimposed for a 60 second sample containing multiple speakers. The bin count has been scaled to a maximum of 1. Compared to digital QAM constellations (which would appear as Dirac functions on this plot), the RADE constellation looks like noise. Plotting individual elements of $\mathbf{z}$ produces a similar histogram (no obvious structure).
@@ 2 RADE Design
    We perform acquisition, synchronisation, and sample rate conversion using well-known DSP techniques. Alternatively, these tasks can also be accomplished using ML [13][14].
@@
The OFDM frame is shown in Fig. 4. Pilot symbols are periodically inserted into each OFDM carrier. After the IDFT stage, we insert a cyclic prefix to guard against inter-symbol interference. To achieve an efficient ratio of pilots to data symbols, we place the QAM symbols from three consecutive $\mathbf{z}$ vectors (120 complex QAM symbols) in each OFDM frame, leading to an OFDM frame duration (and algorithmic delay) of 120 ms.
At the receiver the pilots are used to estimate the time-varying phase of the channel (equalisation), and for initial acquisition (coarse frequency, frame sync) of the received signal. Phase equalisation also allows small frequency offsets (±2$\pm$ 2± 2 Hz) to be handled and frequency drift tracked. Neural networks are sensitive to magnitude scaling, so we also use the pilot symbols for coarse magnitude equalisation (gain control) of the received $\mathbf{\hat{z}}$ signal.
The insertion of pilot symbols and the cyclic prefix consume 2 dB of carrier power that would otherwise be available for payload symbols, and require the symbol rate and hence overall RF bandwidth to be increased by 500 Hz to maintain the payload symbol rate. An additional 2-dB penalty is incurred due to non-ideal performance of the least squares estimation algorithm used for phase equalisation at our low SNR operating point. The total cost of the equalization and cyclic prefix insertion is therefore 4 dB. The resulting waveform has an RF bandwidth of approximately 1500 Hz.
@@ 3 Training
which simulates the saturation of a transmitter power amplifier by compressing the magnitude but retaining the phase. We reasoned that a bottleneck applied to the magnitude of the complex time domain signal would encourage the network to maximise the RMS power (and hence minimse the PAPR) given the channel noise and peak power constraint of the bottleneck. Other approaches that include PAPR in the loss function [15] are also possible.
To train we need to apply the $\mathrm{ctanh}(x)$ bottleneck in the time domain, and simultaneously apply a multipath channel model. Applying the multipath model in the time domain introduces phase rotations and inter-symbol interference (ISI), which would then require equalisation and removal inside the training loop. While possible this would significantly slow down the training process, and require training at the higher rate $F_{s}$ sample rate ($F_{s}/R_{s}=160$ in our design).
@@ 4 Evaluation and Results
Informal listening tests on simulated and over the air samples suggest RADE significantly outperforms SSB. Since intelligibility – more than quality – is the primary goal for HF radio, we use Automatic Speech Recognition (ASR) to evaluate the performance of the proposed system. Five hundred samples from the Librispeech dataset were passed through RADE, SSB, and FreeDV 700D simulations at a range of SNRs, then post processed by the Whisper ASR system [17], and the Word Error Rate (WER) measured (Fig. 6). SSB was simulated by band limiting the input speech to 300-2700 Hz, and applying Hilbert compression such that the mean PAPR was around 8 dB. FreeDV 700D is an open-source HF digital voice protocol using an OFDM modem, rate 1/21/21 / 2 FEC, and the Codec 2 classical vocoder [18]. The MPP channel was simulated using the time domain Watterson model Eq. (4). The Librispeech speech and Watterson model $G_{1}$ and $G_{2}$ datasets used for the evaluation were not part of the training dataset.
We consider two thresholds (a) Link closure - the point where barely intelligible speech can be sent over the system and (b) Good intelligibility, effortless communication. At the 30% WER level (link closure), the ASR results indicate a 4 dB improvement for RADE over SSB for both AWGN and MPP channels. At the 5% WER level (good quality), the improvement is 13 dB for both channels. FreeDV 700D exhibits low speech quality with the Librispeech dataset, although we note it is in common use on HF by trained operators. Similarly, skilled operators can use SSB down to 0 dB SNR. Note the sharp knee in the FreeDV 700D AWGN curve at 2-2- 2 dB SNR, common in digital speech system due to the abrupt breakdown of the FEC. RADE and SSB have a more desirable gradual trade off between SNR and speech quality.
@@
The WER of FARGAN with clean features is very close to that of clean speech, thus our use of classical features and choice of vocoder is not a limiting factor.
@@
Figure 6: Word error rate % versus SNR for simulated AWGN and MPP channels, with clean and FARGAN with clean features as controls. The label SNR3k denotes all signals normalised to a 3000 Hz noise bandwidth.
@@ 4.1 Complexity
The encoder and decoder each require 1 MB of read-only storage (1M weights with 8-bit uniform quantization included in training to facilitate SIMD support) and require 32 MMACs for real-time operation. Both easily operate in real time on a laptop using an unoptimised PyTorch implementation. The overall complexity of the complete system is dominated by the FARGAN vocoder’s 300 MMACs [9].
@@ 4.2 Over the Air Demonstration
To test RADE over real world HF channels licensed Amateur Radio operators from around the world were invited to record a 10 second input sample of their own voice. This was converted to a RADE waveform sample, and concatenated with a Hilbert-compressed version of the same input sample to emulate an analog SSB signal with a 6-8 dB PAPR. Participants then played the concatenated sample through their SSB transmitters over real world HF Radio channels to remote KiwiSDR receivers of their choice (KiwiSDRs are SSB radio receivers that are connected to the public Internet). The received off air signal was processed to obtain a file of SSB and RADE output speech samples, and an estimate of channel SNR. The use of stored files enabled the same Tx signal to be transmitted to different KiwiSDR receivers under different channel conditions, and different transmit power levels. Tests were performed in English, Japanese, Cantonese, and German, over a variety of HF channels of up to 14,000 km (e.g. direct transmission from North America to Australia). Speech samples are available at [19]. Source code is available on GitHub [20] in the waspaa_2025 branch.
@@ 5 Conclusion
We have combined a ML vocoder, ML autoencoder and classical DSP OFDM to build a system capable of sending speech over HF radio channels. It is robust to AWGN and multipath channel impairments, and the transmit signal has a PAPR of less than 1 dB. Unusually for HF speech systems, the audio bandwidth is 8000 Hz, despite requiring just 1500 Hz of RF bandwidth. Our ASR simulation and real world demonstration show performance significantly exceeding that of analog SSB at the same SNR. Unlike classical DSP HF systems, speech quality improves gradually with channel SNR without any mode switching. Interestingly, our system also shows robustness to channel impairments we did not train for, e.g. impulse noise.